how do you graph 4x-3y=12
Answer by mananth(16946)

4x-3y=12
Take the values of -3 , 0 and 3 for x
when x=-3,
-12-3y=12
-3y=12+12
-3y=24
y=-8
when x=0
0-3y=12
y=-12/3
y=-4
when x=3
12-3y=12
-3y=12-12
-3y=0
y=0

Plot the line for co-ordinates (-3 -8) ( 0 -4) (3 0)